# CERN-TH-2025-027

### v1

@@ Abstract
We show that in realistic models where primordial black holes are formed due to the collapse of sizeable inflationary perturbations, their initial spatial clustering beyond Poisson distribution does not play any role in the binary mergers, including sub-solar primordial black holes, responsible for the gravitational waves detectable by LIGO-Virgo-KAGRA. This is a consequence of the existing FIRAS CMB distortion constraints on the relevant scales. This conclusion might not hold for lighter masses potentially accessible by future gravitational wave observations.
@@ 1 Introduction
The physics of Primordial Black Holes (PBHs) has garnered significant interest in recent years (see [1, 2, 3, 4, 5] for some reviews), largely due to the numerous detections of gravitational waves (GWs) originating from BH binary mergers [6, 7, 8, 9, 10] and the hypothesis that some of these may be of primordial origin [11, 12, 13].
@@ 3 The PBH clustering
where $k_{\text{\tiny min}}$ and $k_{\text{\tiny max}}=k_{\text{\tiny min}}\cdot\Delta$ define the minimum and maximum scales, with $\Delta$ being the dimensionless width of the spectrum. With such a power spectrum, the lightest PBHs form when the smallest scale $\sim k^{-1}_{\text{\tiny max}}$ re-enters the horizon. They dominate the mass function over the heavier PBHs formed later because their density is diluted much slower than radiation [108, 109] (an effect which has not been accounted for in Ref. [110]). As we will show, such light PBHs are clustered up to the scale $\sim k^{-1}_{\text{\tiny min}}$. The integral in Eq. (3.5) can be solved analytically, giving
@@ 5 Conclusions
Discovering the presence of PBHs is one of the main targets of the new era of gravitational wave astronomy. Assessing the merger rate of PBH binaries has become therefore imperative. In this paper, we have made a simple, but relevant remark, that the initial clustering of PBH in the mass range relevant for LVK does not affect the standard estimate of the PBH binary merger rate [149] which assumes an initial PBH Poisson distribution. Our conclusions apply to realistic models in which PBHs are formed via the collapse of sizeable fluctuations generated during inflation. They are also robust in the sense that we have assumed a broad power spectrum for the curvature perturbation, thus maximally enhancing the possible clustering.
PBH may also form in alternative scenarios [158]. For the case of phase transitions (PTs) [159, 160, 161, 162], for example, the correlation length is limited by causality as PTs cannot produce super-Hubble correlations at formation time. The relevant scales that can significantly affect PBH binaries are much larger. So, we do not expect this scenario to avoid our conclusion. We leave the investigation of alternative cases for future work.

### v2

@@ Abstract
We show that in common inflationary models where primordial black holes are formed due to the collapse of sizeable inflationary perturbations, their initial spatial clustering beyond the Poisson distribution does not affect the binary mergers— including sub-solar primordial black holes—responsible for the gravitational waves detectable by LIGO-Virgo-KAGRA. This is a consequence of the existing FIRAS CMB distortion constraints on the relevant scales. However, this conclusion may not hold for lighter masses potentially accessible by future gravitational wave observations and for multi-field inflation scenarios, where the curvature perturbation can acquire a large-scale modulation from a field that is not the primary source of the perturbation.
@@ 1 Introduction
The physics of Primordial Black Holes (PBHs) has garnered significant interest in recent years (see [1, 2, 3, 4, 5] for some reviews), largely due to the numerous detections of gravitational waves (GWs) originating from BH binary mergers [6, 7, 8, 9, 10] and the hypothesis that some of these may be of primordial origin [11, 12, 13, 14, 15].
@@ 3 The PBH clustering
where $k_{\text{\tiny min}}$ and $k_{\text{\tiny max}}=k_{\text{\tiny min}}\cdot\Delta$ define the minimum and maximum scales, with $\Delta$ being the dimensionless width of the spectrum. With such a power spectrum, the lightest PBHs form when the smallest scale $\sim k^{-1}_{\text{\tiny max}}$ re-enters the horizon. They dominate the mass function over the heavier PBHs formed later because their density is diluted much slower than radiation [110, 111]. As we will show, such light PBHs are clustered up to the scale $\sim k^{-1}_{\text{\tiny min}}$. The integral in Eq. (3.5) can be solved analytically, giving
@@ 5 Conclusions
Discovering the presence of PBHs is one of the main targets of the new era of gravitational wave astronomy. Assessing the merger rate of PBH binaries has become therefore imperative. In this paper, we have made a simple, but relevant remark, that the initial clustering of PBH in the mass range relevant for LVK does not affect the standard estimate of the PBH binary merger rate [150] which assumes an initial PBH Poisson distribution. Our conclusions apply to common single-field models in which PBHs are formed via the collapse of sizeable fluctuations generated during inflation. They are also robust in the sense that we have assumed a broad power spectrum for the curvature perturbation, thus maximally enhancing the possible clustering.
Here we stress that our claim does not apply for multi-field inflation scenarios [114, 60, 159], where the curvature perturbation can acquire a large-scale modulation from a field that is not the primary source of the perturbation. Moreover, PBH may also form in alternative scenarios [160]. For the case of phase transitions (PTs) [161, 162, 163, 164], for example, the correlation length is limited by causality as PTs cannot produce super-Hubble correlations at formation time. The relevant scales that can significantly affect PBH binaries are much larger. So, we do not expect this scenario to avoid our conclusion. We leave the investigation of alternative cases for future work.